Circular No. 7113

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

U SCORPII P. Schmeer, Bischmisheim, Germany, reports an outburst of this recur-

rent nova, which last flared in 1987 (cf. IAUC 4395). Confirmatory observations have been received via E. Waagen at the AAVSO. Visual magnitudes: Feb. 25.194 UT, 9.5 (Schmeer); 25.562, 7.6 (L. Shaw, Piñole, CA); 25.625, 8.0 (M. Mattiazzo, Wallaroo, South Australia).

XTE J1550−564

J. Casares, Instituto de Astrofisica de Canarias; G. Dubus, University of Amsterdam; and L. Homer, University of Oxford, report: “We have obtained B, V and R images and medium-resolution spectroscopy (0.38 nm) of the x-ray transient XTE J1550−564 (IAUC 7008) on Feb. 17–18 with the 3.5-m New Technology Telescope at the European Southern Observatory, La Silla. We measured B = 16.6 ± 0.1, V = 16.3 ± 0.1, R = 15.5 ± 0.1 on 1998 Sept. 19.4, so the transient has again reached the optical brightness of the outburst last September (IAUC 7010) during this current x-ray reflare (IAUC 7098, 7013). Our spectra show a reddened continuum dominated by strong interstellar features and broad emission lines of Hα (FWHM = 815 km/s, EW = 0.3 nm), He II 468.6-nm (FWHM = 1150 km/s, EW = 0.2 nm) and the Bowen blend (FWHM = 1370 km/s, EW = 0.3 nm), characteristic of x-ray transients in outburst. We also identify weak emission lines of He I 587.5-nm and 667.8-nm and Hβ. The Hα and He II 468.6-nm lines are double-peaked with a separation of ∼ 380 km/s and exhibit a factor of 2–3 variability in their equivalent widths between the two nights. We measure equivalent widths −3.6 for Na D (588.9–589.5 nm) and −0.30 for 661.3 nm, from which we estimate an approximate E(B−V) = 1.0±0.1. More optical observations of this new blackhole candidate are very much required.”

COMET 52P/HARRINGTON-ABELL Total visual magnitude estimates: Jan. 5.02 UT, 10.8 (J. Bortle,

Stormville, NY, 0.41-m reflector); 10.74, 10.8, (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 16.65, 11.5 (A. Pearce, Nedlands, Western Australia, 0.41-m reflector); 23.95, 10.8 (M. Meyer, Frauenstein, Germany, 0.25-m reflector); 27.58, 11.1 (N. Biver, Oahu, HI, 0.26-m reflector); Feb. 5.79, 11.1 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 9.89, 10.9 (A. Baransky, Kiev, Ukraine, 0.11-m reflector); 13.74, 10.7 (S. Yoshida, Ibaraki, Japan, 0.25-m reflector); 16.91, 11.9 (J. Carvajal, Madrid, Spain, 0.21-m reflector); 22.74, 10.5 (M. Lehký, Hradec Králové, Czech Republic, 25 × 100 binoculars).

1999 February 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Brian G. Marsden